# OneBit: Towards Extremely Low-bit Large Language Models

### v2

@@ 4.2 Main Results
Table 2 compares our method with other typical strong baselines on different models. Due to space limitations, results of LLaMA2-7B/13B are listed in Appendix A.3. In various model sizes, our 1-bit weight quantization method obviously outperforms others under the W2A16 setting. Moreover, the effectiveness of QAT based methods consistently improves as the model size increases, whereas the result of the PTQ method, GPTQ, may degrade when model size increases (e.g., from 7B to 13B on LLaMA). This demonstrates that QAT-based method can achieve stable results in extremely low-bit quantization. Specifically, our method approaches the performance of FP16 more closely as the model size increases. For instance, when scaling from LLaMA-7B to LLaMA-13B, the perplexity of the FP16 model decreases by only 0.59, whereas our method sees a reduction of 1.20.
@@
For zero-shot accuracy, although all methods inevitably have some degradation, our method achieves the closest performance to the FP16 baseline among most models. On the OPT-1.3B/2.7B model, our method shows smaller performance loss on most tasks such as PIQA and ARC-e. Additionally, the loss of other tasks is negligible compared with the second-best baseline, OmniQuant. On the LLaMA-7B model, our method notably outperforms OmniQuant in all tasks except ARC-e/ARC-c, averaging about a 4% improvement overall.
@@ 4.3 Problem Solving Ability
We have demonstrated the superior performance of our method under the W1A16 setting, compared to other representative baselines. Although all methods inevitably face performance degradation in 1-bit weight quantization, it remains of interest how our method fares in solving practical problems among the various approaches to reducing model size. For instance, directly training smaller models (Zhang et al., 2024) or employing low-rank decomposition to reduce the number of parameters.
To this end, we consider two crucial abilities of LLMs: commonsense reasoning and world knowledge. For commonsense reasoning, we use the 6 tasks (Hellaswag, etc.) and settings described in Section 4.2. For world knowledge, we examine it using the Massive Multi-task Language Understanding (MMLU; Hendrycks et al., 2021), a benchmark that covers a wide range of domains and knowledge. We compare the following 4 models:
Pythia-1.0B Biderman et al. (2023). A well-trained model released by EleutherAI whose memory footprint is 1.54x that of our 7B model.
TinyLLaMA-1.1B (Zhang et al., 2024). A model with the same structure as the LLaMA models, which undergoes continued training. To compare fairly, we use the checkpoint at 10k training steps, which is 2x that of our model.
LowRank LLaMA (Noach and Goldberg, 2020). Decompose every weight matrix in Linear layers to two low-rank matrices and learn from the original LLaMA-7B model by KD in the same setting of OneBit-7B.
@@ 5.1 Efficiency
Table 3: Compression ratio of LLaMA models.
Figure 4: Tradeoff between model size and perplexity.
Figure 5: Training process of OneBit-7B.
@@ 5.2 Robustness
Existing work (Wang et al., 2023) has already noted the instability within QAT. Extremely low-bit quantization makes the training process highly sensitive to the learning rate, making it difficult for the model to converge when the rate is too small or too large. This is primarily due to the large magnitude of gradients generated as the weight elements fluctuate between +1 and -1, leading to substantial fluctuations in the output of Linear layers. Experiments demonstrate that OneBit shows more stable training process and is not sensitive to learning rates. Please refer to Appendix A.5 for more details.
@@ 5.3 Effect of Different Components
We discover that models might experience floating-point overflow during the QAT process. As depth increases, the activation can become progressively larger. We tackle it using Post-LayerNorm instead of Pre-LayerNorm. In contrast, Pre-LayerNorm may occasionally be ineffective.
@@
The main structural difference between OneBit and BitNet (Wang et al., 2023) is the two value vectors, which are demonstrated to be effective in Section 4.2. Please refer to Appendix A.5 for more details of comparison.
@@ 6 Conclusion
We propose a model structure for 1-bit weight quantization and a corresponding parameter initialization method. Extensive experiments on models of various sizes and series demonstrate that OneBit has clear advantages over representative strong baselines and achieves a good tradeoff between model size and performance. We further analyze the capabilities of such extremely low-bit quantized models and provide guidance for future research.
Limitation
Although our proposed method significantly reduces the memory footprint of LLMs, bringing hope for efficient deployment of them, there are still some limitations. Firstly, compared to the original model, our extremely low-bit quantization inevitably incurs a performance loss. Additionally, we are yet to understand the mathematical principles behind the optimal parameters of the 1-bit quantized model, thus capability transfer can only be achieved through the costly process of KD. Fortunately, this cost is a one-time expense. Moreover, due to the unique nature of 1-bit quantization, our method can not be naturally extended to higher bit-width. Lastly, we have not considered the activation quantization and leave it as future work.
@@ A.1 Proofs of Propositions
Wiki2 C4 Winogrande Hellaswag PIQA BoolQ ARC-e ARC-c Avg.
@@
Table 4: Results of LLaMA2. We bold the best scores.
A.2 Details on Baselines
  • GPTQ (Frantar et al., 2022): We employ the open-source code released by the author. Both OPT models and LLaMA models take 128 2048-token samples from the C4 dataset to calibrate the quantized model. For LLaMA models, we apply the activation order heuristic according to the recommendation from the code.
  • LLM-QAT (Liu et al., 2023): We reimplement this method to adapt the W2A16 setting, as LLM-QAT is not designed for 2-bit weight quantization. We also do not quantize the KV Cache. When quantizing the weight matrix in Linear layer, we use symmetric MinMax quantization in which the zero-point is set to 0. The training hyper-parameters are the same as ours. Please refer to the training details in Section 4.1.
  • OmniQuant (Shao et al., 2023): We employ the open-source code released by the author. Both OPT models and LLaMA models take 128 2048-token samples from the WikiText2 dataset to calibrate the quantized model. The learning rate for learnable weight clipping and equivalent transformation is set to 5e-3 and 1e-2, respectively. We use a batch size of 1 and train 40 epochs for each model. For OPT models, both learnable weight clipping and equivalent transformation are leveraged. For LLaMA models, only learnable weight clipping is used.
A.3 Results of LLaMA2
A.4 Instrution Following Ability
Instruction following is an important ability of LLMs (Radford et al., 2019; Brown et al., 2020; Peng et al., 2023). Beyond the discussion on model abilities and efficiency before, we also focus on the instruction following ability of extremely low-bit models, which is closely related to their practical usability. In this subsection, we empirically study this capability of our quantized model. We fine-tune the model for 3 epochs using the alpaca_en_52k dataset and alpaca templates (Taori et al., 2023), then observe the generation in both zero-shot and few-shot settings before and after fine-tuning. During training, the learning rate is set to 1e-7 and the batch size to 32. Other parameters are consistent with Section 4.1.
@@ A.4 Instrution Following Ability
D (Correct answer)
@@ A.5 Comparison with BitNet
Recently, BitNet (Wang et al., 2023) introduces a 1-bit model architecture and applies the architecture to train models from scratch, demonstrating the feasibility and application value of the 1-bit model structure. In this paper, we attempt to combine 1-bit quantization with knowledge distillation to quantize the LLaMA-7B model. Unfortunately, despite following the suggestion to use larger learning rates, the behavior remains unstable during training.
@@
A.6 Average Bit-width of Linear Layer

### v3

@@ 4.2 Main Results
Table 2 compares our method with other typical strong baselines on different models. Due to space limitations, results of LLaMA2-7B/13B are listed in Appendix A.3. In various model sizes, our 1-bit weight quantization method obviously outperforms others under the W2A16 setting. Moreover, the effectiveness of QAT based methods consistently improves as the model size increases, whereas the result of the PTQ method, GPTQ, may degrade when model size increases (e.g., from 7B to 13B on LLaMA). This demonstrates that QAT-based method can achieve stable results in extremely low-bit quantization. Specifically, our method approaches the performance of FP16 more closely as the model size increases. For instance, when scaling from LLaMA-7B to LLaMA-13B, the perplexity (on C4) of the FP16 model decreases by only 0.47, whereas our method sees a reduction of 1.15.
@@
For zero-shot accuracy, although all methods inevitably have some degradation, our method achieves the closest performance to the FP16 baseline among most models. On the OPT-1.3B/2.7B model, our method shows smaller performance loss on most tasks such as PIQA and ARC-e. Additionally, the loss of other tasks is negligible compared with the second-best baseline, OmniQuant. On the LLaMA-7B model, our method also notably outperforms OmniQuant in most tasks except BoolQ/ARC-e, averaging about a 4% improvement overall.
@@ 4.3 Problem Solving Ability
We have demonstrated the superior performance of our method under the W1A16 setting, compared to other representative baselines. Although all methods inevitably face performance degradation in 1-bit weight quantization, it remains of interest how our method fares in solving practical problems among the various approaches to reducing model size. For instance, directly training smaller models [42] or employing low-rank decomposition to reduce the number of parameters.
To this end, we consider two crucial abilities of LLMs: commonsense reasoning and world knowledge. For commonsense reasoning, we use the 6 tasks (Hellaswag, etc.) and settings described in Section 4.2. For world knowledge, we examine it using the Massive Multi-task Language Understanding (MMLU; 15), a benchmark that covers wide domains and knowledge. We compare the following 4 models:
Pythia-1.0B [3] A well-trained model released by EleutherAI whose memory footprint is 1.54x that of our OneBit-7B model.
TinyLLaMA-1.1B [42] A model with the same structure as the LLaMA models, which undergoes continued training. To compare fairly, we use the checkpoint at 10k training steps, which is 2x that of our OneBit-7B model.
LowRank LLaMA [24] Decompose every weight matrix in Linear layers to two low-rank matrices and learn from the original LLaMA-7B model by KD in the same setting of OneBit-7B.
@@ 5.1 Efficiency
Figure 4: Tradeoff between size and PPL.
Figure 5: Training process of OneBit-7B.
Table 3: Compression ratio of LLaMA models.
@@ 5.2 Robustness
Existing work [38] has already noted the instability within QAT. Extremely low-bit quantization makes the training process highly sensitive to the learning rate, making it difficult for the model to converge when the rate is too small or too large. This is primarily due to the large magnitude of gradients generated as the weight elements fluctuate between +1 and -1, leading to substantial fluctuations in the output of Linear layers. Experiments demonstrate that OneBit shows more stable training process and is not sensitive to learning rates. Please refer to Appendix A.5 for more details.
@@ 5.3 Effect of Different Components
We discover that there might be floating-point overflow during the QAT process. As depth increases, the activation can become progressively larger. We tackle it using Post-LayerNorm instead of Pre-LayerNorm. In contrast, Pre-LayerNorm may occasionally be ineffective.
@@
The main structural difference between OneBit and BitNet [38] is the two value vectors, which are demonstrated to be effective in Section 4.2. They facilitate stable training and the knowledge transfer process. Please refer to Appendix A.5 for more details of comparison.
@@ 6 Conclusion
We propose a novel model structure for 1-bit weight quantization and a corresponding parameter initialization method to address the difficulty in 1-bit quantization. Extensive experiments on LLMs of various sizes and series demonstrate that OneBit has clear advantages over representative strong baselines and achieves a good tradeoff between model size and performance. We further analyze the capabilities of such extremely low-bit quantized models and provide guidance for future research.
Limitations
Although our proposed method significantly reduces the memory footprint of LLMs, bringing hope for efficient deployment of them, there are still some limitations. Firstly, compared to the original model, our extremely low-bit quantization inevitably incurs a performance loss. Additionally, we are yet to understand the mathematical principles behind the optimal parameters of the 1-bit quantized model, thus capability transfer can only be achieved through the relatively costly process of KD. Fortunately, this cost is a one-time expense. Moreover, due to the unique nature of 1-bit quantization, our method can not be naturally extended to higher bit-width. Lastly, we have not considered the activation quantization and leave it as future work.
@@ A.1 Proofs of Propositions
A.2 Details on Baselines
  • GPTQ [14]: We employ the open-source code released by the author. Both OPT models and LLaMA models take 128 2048-token samples from the C4 dataset to calibrate the quantized model. For LLaMA models, we apply the activation order heuristic according to the recommendation from the code.
  • LLM-QAT [21]: We reimplement this method to adapt the W2A16 setting, as LLM-QAT is not designed for 2-bit weight quantization. We also do not quantize the KV Cache. When quantizing the weight matrix in Linear layer, we use symmetric MinMax quantization in which the zero-point is set to 0. The training hyper-parameters are the same as ours. Please refer to the training details in Section 4.1.
  • OmniQuant [30]: We employ the open-source code released by the author. Both OPT models and LLaMA models take 128 2048-token samples from the WikiText2 dataset to calibrate the quantized model. The learning rate for learnable weight clipping and equivalent transformation is set to 5e-3 and 1e-2, respectively. We use a batch size of 1 and train 40 epochs for each model. For OPT models, both learnable weight clipping and equivalent transformation are leveraged. For LLaMA models, only learnable weight clipping is used.
A.3 Results of LLaMA2
Table 4: Results of LLaMA2. We bold the best scores.
@@ A.3 Results of LLaMA2
Wiki2 C4 Wino. Hella. PIQA BoolQ ARC-e ARC-c Avg.
@@
D
A.4 Instrution Following Ability
Instruction following is an important ability of LLMs [27, 5, 26]. Beyond the discussion on model abilities and efficiency before, we also focus on the instruction following ability of extremely low-bit models, which is closely related to their practical usability. In this subsection, we empirically study this capability of our quantized model. We fine-tune the model for 3 epochs using the alpaca_en_52k dataset and alpaca templates [33], then observe the generation in both zero-shot and few-shot settings before and after fine-tuning. During training, the learning rate is set to 1e-7 and the batch size to 32. Other parameters are consistent with Section 4.1.
@@ A.5 Comparison with BitNet
Recently, BitNet [38] introduces a 1-bit model architecture and applies the architecture to train models from scratch, demonstrating the feasibility and application value of the 1-bit model structure. In this paper, we attempt to combine 1-bit quantization with knowledge distillation to quantize the LLaMA-7B model. Unfortunately, despite following the suggestion to use larger learning rates, the behavior remains unstable during training.
@@
Table 6: Ablation study of different loss on LLaMA-7B. “ATTN” means attention score alignment.
Loss Setting Perplexity()($\downarrow$)( ↓ ) Zero-shot Accuracy()($\uparrow$)( ↑ )
Wiki2 C4 Wino. Hella. PIQA BoolQ ARC-e ARC-c Avg.
$\mathcal{L}_{\mathrm{KD}}$ 13.48 14.57 50.83 35.14 62.89 60.46 37.33 26.37 45.50
$\mathcal{L}_{\mathrm{KD}}+\mathcal{L}_{\mathrm{MSE}}$ ($\alpha=1$) 10.19 11.40 58.48 51.54 68.01 57.28 42.47 30.20 51.33
$\mathcal{L}_{\mathrm{KD}}+\mathcal{L}_{\mathrm{MSE}}$ ($\alpha=10$) 10.38 11.56 60.30 50.73 67.46 62.51 41.71 29.61 52.05
$\mathcal{L}_{\mathrm{KD}}+\mathcal{L}_{\mathrm{MSE}}+\mathcal{L}_{\mathrm{ATTN}}$ NAN NAN - - - - - - -
A.6 Discussion on Knowledge Distillation
Although knowledge distillation is not the main contribution of this paper, we nevertheless provide the rationale behind certain settings used in our experiments to explain the necessity of these configurations.
We firstly explain the role of different loss functions in guiding the process of knowledge transfer. Fundamentally, distillation loss alone can achieve a satisfactory transfer process (comparing to other baselines). Additionally, as shown in the Table 6, aligning the hidden states between layers can result in a quantized model with better perplexity. However, further incorporating attention score alignment on this basis leads to the model failing to converge. LLM-QAT [21] has conducted similar experiments on quantization-aware knowledge distillation loss and concluded that using only the distillation loss yields the best results. The difference in conclusions may stem from two factors. On one hand, due to our adoption of a novel model architecture, which differs from theirs, the optimal usage of loss functions may be different as well. On the other hand, as we focus on extremely low bit-width compression, each layer of the model suffers significant information loss compared to the teacher model. The regularization of hidden states between layers may help reduce the variance in the learning process, thus demonstrating stronger generalization.
Furthermore, we also discuss the cost of our quantization method. Using LLaMA-7B as an example, quantizing the model with our method requires approximately 7 days on 8 A100-80GB GPUs. In comparison, training the LLaMA-7B model from scratch consumes 82,432 GPU hours [34]. The quantization time, being less than 2% of the pretraining time, is still an acceptable cost.
A.7 Average Bit-width of Linear Layer